# Better than classical? The subtle art of benchmarking quantum machine learning models

### 11 Mar 2024 v1

@@ II.2 Summary of the models
The 12 models we implemented can be split into three families: quantum neural networks, quantum kernel methods and quantum convolutional neural networks. Here we explain the principle of each family and give a brief overview of the central idea of each model. Technical terms commonly known within the quantum machine learning community, but not necessarily clear to every reader, are highlighted in orange italic font and explained in a glossary at the end of this paper (see Appendix A). More detailed descriptions of the models can be found in Appendix B, and Table 1 provides the basic facts at a glance. We use camel case to refer to models, and the names correspond to the classes used in the software package.
@@ II.2.3 Quantum convolutional neural network models
Table 1: Overview of models used in the benchmarks. For definitions of the terms, consult the glossary in Appendix A. More details on the models are found in Appendix B.
@@ II.3 Implementation
    Faithful implementation: We carefully deduced the model design and training procedure from the paper. If specific details needed for implementation were missing in the paper, they were defined based on what appeared natural given our own judgement (see also Appendix B).
@@
    Convergence criteria: All quantum neural network models and quantum convolutional models used the same convergence criterion, which was based on tracking the variance of the loss values over time and stopping when the mean loss value did not change significantly over 200 parameter updates [^10] As with all variational models, it can be difficult to know whether a model has converged or if the model is stuck on a particularly flat part of the optimisation landscape, and test accuracies can both improve or worsen with more training. The choice to decide on convergence over 200 updates is therefore to some extent arbitrary. (see Appendix C for details). Loss histories were routinely inspected by eye to ensure the criterion was working as desired. Quantum kernel methods followed the default convergence criterion of the specific scikit-learn model class (SVC, Logistic Regression) that they employ. Pure scikit-learn baseline models (MLPClassifier, SVC) used their default convergence criterion except for the max_iter parameter of MLPClassifier which was increased to 3000. If a training run did not converge during grid search, which happened very rarely, that particular result was ignored in the hyperparameter optimisation procedure.
@@ III Datasets
Figure 5: Illustrative examples of datasets created by the different data generation procedures. For the scatter plots, the two classes are shown in blue and orange, and training points are shown in round vs. test points in an ‘x’ shape. The linearly separable panel shows data for the linearly separable benchmark in 2 and 3 dimensions. The left two plots for the MNIST data correspond to 2d and 3d mnist pca data, and the rightmost image shows examples from the mnist-cg dataset for 32 x 32, 16 x 16, 8 x 8 and 4 x 4 pixel grids. The hidden manifold examples correspond to a $1$d (left) and $2$d (center) and $3$d (right) manifold embedded into $3$ dimensions. The bars and stripes panel shows examples from the bars & stripes dataset for a 16 x 16 pixel grid. The examples from the two curves diff benchmark show a degree of $2,10,20$ for the Fourier series, embedding the curves into $10$ dimensions (of which three are plotted). The hyperplanes panel shows data from the hyperplanes diff benchmark, where there are two (left) and five (right) hyperplanes used to decide the class labels.
@@
With this in mind we define $6$ data generation procedures, with which we generate data for $10$ benchmarks (in the following named in capital letters). Each benchmark consists in turn of several datasets that differ by varying parameters in the data generation procedure (in most cases the input dimension). Overall, the benchmarks consist of $160$ individual datasets. While the $6$ data generation procedures and their associated benchmarks are summarised in the following list and illustrated in Figure 5, the precise generation settings can be found in Appendix D.
@@ IV Hyperparameter tuning
We conduct an extensive hyperparameter search for all models and datasets in all our experiments, using a full grid search algorithm implemented by the scikit-learn package [71] with the default five-fold cross-validation, using the accuracy score [^15] The accuracy can be an overly simplistic measure for some classification tasks and especially with unbalanced data [88]. However, we utilise it here for its clarity in interpretation, and since our datasets are balanced. to pick the best model. While there are more sophisticated techniques [89, 90], a full grid search has the advantage of simplicity, and allows us to extract and study unbiased information about the hyperparameter landscapes. We remark that the number of hyperparameters varies significantly from model to model since our aim was to follow the proposals in the original papers. In some cases, we were forced to select a subset due to the exponential increase in grid size with the number of hyperparameters. In these cases, the most relevant hyperparameters were selected by analysing preliminary results first on smaller-scale data. We describe the hyperparameter ranges for each model in Appendix B and summarise a few findings of potential interest here.
@@ V.1 Out-of-the-box classical models outperform quantum models
Breaking down the fully aggregated results, one finds that the rankings of models between benchmarks were surprisingly consistent (see Appendix E for a full report). For example, in four out of seven benchmarks used for QNN models, the MLPClassifier ranked first with the DressedQuantumCircuitClassifier coming second, whereas in the remaining three benchmarks the roles between these (conceptually very similar) models were reversed. The DataReuploadingClassifier and QuantumBoltzmannMachine (which was not run on the 10-d DIFF benchmarks) usually share places three and four. In five out of the seven benchmarks, the bottom ranks are taken by the CircuitCentricClassifier and TreeTensorClassifier as the worst-performing models – interestingly the only two models based on amplitude encoding.
@@ V.2 No systematic overfitting or scaling effects
There are three interesting general observations of effects one might expect but which we did not observe. These are exemplified with the selected detailed benchmark results shown in Figure 11. Firstly, although not all quantum models – in particular the QNN methods – use explicit regularisers, they do not show systematic overfitting of the training data. (Some exceptions can be seen in the hidden manifold, hidden manifold diff and two curves diff benchmarks shown in Appendix E, for which also the classical models struggle with overfitting.)
Secondly, we do not observe any improvement in performance of the quantum models relative to the classical baselines for increasing problem difficulties. For the difficulty-controlled benchmarks, hyperplanes diff (Figure 11) and hidden manifold diff (Appendix E), the trends of the quantum models’ test accuracies generally follow the trend of the classical model. For the difficulty-controlled benchmark two curves diff (Appendix E), the quantum models perform worse than the corresponding classical method as the difficulty is increased. Interestingly, for the hardest datasets from this benchmark, the classical baseline models achieve high ($>90\%$) test accuracy whereas all quantum models appear to struggle. This is somewhat surprising, since the embedded curve that defines the structure of the data is a Fourier series, and one may expect quantum models to have a natural bias for this kind of data [91].
@@ V.3 Quantum circuits without entanglement do well
Replotting the test accuracies from Figure 11 with the new models added we see in Figure 12 that the non-entangled models do surprisingly well. This can be confirmed by including the separable models in the ranking results from Figure 10 (see Appendix E). One finds that compared to the quantum kernel methods, the SeparableKernelClassifier takes second rank after the SVC. Among the QNNs, the SeparableVariationalClassifier is only consistently beaten by the MLPClassifier, DressedQuantumCircuitClassifier and QuantumBoltzmannMachine, as well as occasionally by DataReuploadingClassifier.
@@ VI.2 What makes data reuploading work?
While the term “data reuploading” is often used rather generally to describe the layer-wise repetition of an encoding, there are a few other distinctive features in the original model we implemented here. For example, the inputs are individually rescaled by classical trainable parameters before feeding them into quantum gates, which in the picture of quantum models as a Fourier series [91] re-scales the frequency spectrum that the Fourier series is built from. Furthermore, there is no separation between the embedding and variational part of the circuit; instead the embedding layer is trainable (see Eq. 14 in App. B), leading to a sort of “trainable data encoding” (a feature that the QuantumMetricLearner also exhibits). Furthermore, the cost function differs from standard approaches as it measures the fidelity to certain states associated with a class label, and contains more trainable parameters. Which of these features is important for the success of the model - or is it a combination of them?
@@ Acknowledgements.
Appendix A Glossary of quantum machine learning concepts
@@
Appendix B Detailed description of the models
@@ B.18 SeparableKernelClassifier
Appendix C Convergence criteria of variational models
@@
Appendix D Details on the datasets used in this study
@@ D.3 Downscaled MNIST
D.4 Hidden manifold model
@@ D.6 Hyperplanes and parity
Appendix E Collection of detailed results

### 14 Mar 2024 v2

@@ II.2 Summary of the models
The 12 models we implemented can be split into three families: quantum neural networks, quantum kernel methods and quantum convolutional neural networks. Here we explain the principle of each family and give a brief overview of the central idea of each model. Technical terms commonly known within the quantum machine learning community, but not necessarily clear to every reader, are highlighted in orange italic font and explained in a glossary at the end of this paper (see Appendix B). More detailed descriptions of the models can be found in Appendix C, and Table 1 provides the basic facts at a glance. We use camel case to refer to models, and the names correspond to the classes used in the software package.
@@ II.2.3 Quantum convolutional neural network models
Table 1: Overview of models used in the benchmarks. For definitions of the terms, consult the glossary in Appendix B. More details on the models are found in Appendix C.
@@ II.3 Implementation
    Faithful implementation: We carefully deduced the model design and training procedure from the paper. If specific details needed for implementation were missing in the paper, they were defined based on what appeared natural given our own judgement (see also Appendix C).
@@
    Convergence criteria: All quantum neural network models and quantum convolutional models used the same convergence criterion, which was based on tracking the variance of the loss values over time and stopping when the mean loss value did not change significantly over 200 parameter updates [^10] As with all variational models, it can be difficult to know whether a model has converged or if the model is stuck on a particularly flat part of the optimisation landscape, and test accuracies can both improve or worsen with more training. The choice to decide on convergence over 200 updates is therefore to some extent arbitrary. (see Appendix D for details). Loss histories were routinely inspected by eye to ensure the criterion was working as desired. Quantum kernel methods followed the default convergence criterion of the specific scikit-learn model class (SVC, Logistic Regression) that they employ. Pure scikit-learn baseline models (MLPClassifier, SVC) used their default convergence criterion except for the max_iter parameter of MLPClassifier which was increased to 3000. If a training run did not converge during grid search, which happened very rarely, that particular result was ignored in the hyperparameter optimisation procedure.
@@ III Datasets
Figure 5: Illustrative examples of datasets created by the different data generation procedures. For the scatter plots, the two classes are shown in blue and orange, and training points are shown in round vs. test points in an ‘x’ shape. The linearly separable pannel shows data for the linearly separable benchmark in 2 and 3 dimensions. The left two plots for the MNIST data correspond to 2d and 3d mnist pca data, and the rightmost image shows examples from the mnist-cg dataset for 32 x 32, 16 x 16, 8 x 8 and 4 x 4 pixel grids. The hidden manifold examples correspond to a $1$d (left) and $2$d (center) and $3$d (right) manifold embedded into $3$ dimensions. The bars and stripes panel shows examples from the bars & stripes dataset for a 16 x 16 pixel grid. The examples from the two curves diff benchmark show a degree of $2,10,20$ for the Fourier series, embedding the curves into $10$ dimensions (of which three are plotted). The hyperplanes pannel shows data from the hyperplanes diff benchmark, where there are two (left) and five (right) hyperplanes used to decide the class labels.
@@
With this in mind we define $6$ data generation procedures, with which we generate data for $10$ benchmarks (in the following named in capital letters). Each benchmark consists in turn of several datasets that differ by varying parameters in the data generation procedure (in most cases the input dimension). Overall, the benchmarks consist of $160$ individual datasets. While the $6$ data generation procedures and their associated benchmarks are summarised in the following list and illustrated in Figure 5, the precise generation settings can be found in Appendix E.
@@ IV Hyperparameter tuning
We conduct an extensive hyperparameter search for all models and datasets in all our experiments, using a full grid search algorithm implemented by the scikit-learn package [71] with the default five-fold cross-validation, using the accuracy score [^15] The accuracy can be an overly simplistic measure for some classification tasks and especially with unbalanced data [88]. However, we utilise it here for its clarity in interpretation, and since our datasets are balanced. to pick the best model. While there are more sophisticated techniques [89, 90], a full grid search has the advantage of simplicity, and allows us to extract and study unbiased information about the hyperparameter landscapes. We remark that the number of hyperparameters varies significantly from model to model since our aim was to follow the proposals in the original papers. In some cases, we were forced to select a subset due to the exponential increase in grid size with the number of hyperparameters. In these cases, the most relevant hyperparameters were selected by analysing preliminary results first on smaller-scale data. We describe the hyperparameter ranges for each model in Appendix C and summarise a few findings of potential interest here.
@@ V.1 Out-of-the-box classical models outperform quantum models
Breaking down the fully aggregated results, one finds that the rankings of models between benchmarks were surprisingly consistent (see Appendix F for a full report). For example, in four out of seven benchmarks used for QNN models, the MLPClassifier ranked first with the DressedQuantumCircuitClassifier coming second, whereas in the remaining three benchmarks the roles between these (conceptually very similar) models were reversed. The DataReuploadingClassifier and QuantumBoltzmannMachine (which was not run on the 10-d DIFF benchmarks) usually share places three and four. In five out of the seven benchmarks, the bottom ranks are taken by the CircuitCentricClassifier and TreeTensorClassifier as the worst-performing models – interestingly the only two models based on amplitude encoding.
@@ V.2 No systematic overfitting or scaling effects
There are three interesting general observations of effects one might expect but which we did not observe. These are exemplified with the selected detailed benchmark results shown in Figure 11. Firstly, although not all quantum models – in particular the QNN methods – use explicit regularisers, they do not show systematic overfitting of the training data. (Some exceptions can be seen in the hidden manifold, hidden manifold diff and two curves diff benchmarks shown in Appendix F, for which also the classical models struggle with overfitting.)
Secondly, we do not observe any improvement in performance of the quantum models relative to the classical baselines for increasing problem difficulties. For the difficulty-controlled benchmarks, hyperplanes diff (Figure 11) and hidden manifold diff (Appendix F), the trends of the quantum models’ test accuracies generally follow the trend of the classical model. For the difficulty-controlled benchmark two curves diff (Appendix F), the quantum models perform worse than the corresponding classical method as the difficulty is increased. Interestingly, for the hardest datasets from this benchmark, the classical baseline models achieve high ($>90\%$) test accuracy whereas all quantum models appear to struggle. This is somewhat surprising, since the embedded curve that defines the structure of the data is a Fourier series, and one may expect quantum models to have a natural bias for this kind of data [91].
@@ V.3 Quantum circuits without entanglement do well
Replotting the test accuracies from Figure 11 with the new models added we see in Figure 12 that the non-entangled models do surprisingly well. This can be confirmed by including the separable models in the ranking results from Figure 10 (see Appendix F). One finds that compared to the quantum kernel methods, the SeparableKernelClassifier takes second rank after the SVC. Among the QNNs, the SeparableVariationalClassifier is only consistently beaten by the MLPClassifier, DressedQuantumCircuitClassifier and QuantumBoltzmannMachine, as well as occasionally by DataReuploadingClassifier.
@@ VI.2 What makes data reuploading work?
While the term “data reuploading” is often used rather generally to describe the layer-wise repetition of an encoding, there are a few other distinctive features in the original model we implemented here. For example, the inputs are individually rescaled by classical trainable parameters before feeding them into quantum gates, which in the picture of quantum models as a Fourier series [91] re-scales the frequency spectrum that the Fourier series is built from. Furthermore, there is no separation between the embedding and variational part of the circuit; instead the embedding layer is trainable (see Eq. 14 in App. C), leading to a sort of “trainable data encoding” (a feature that the QuantumMetricLearner also exhibits). Furthermore, the cost function differs from standard approaches as it measures the fidelity to certain states associated with a class label, and contains more trainable parameters. Which of these features is important for the success of the model - or is it a combination of them?
@@ Acknowledgements.
Appendix A CO2 Emission Table
The following calculations are made using energy consumption data for the Cedar supercomputing cluster at the SciNet HPC Consortium, on which the vast majority of our simulations were run.
Numerical simulations
Total CPU usage [core years] $\approx 27$
Cluster energy consumption (per core) [W] $\approx 11$
Total Energy Consumption Simulations [kWh] $\approx 2600$
[…]
Transport
[…]
Were The Emissions Offset? No
Appendix B Glossary of quantum machine learning concepts
@@
Appendix C Detailed description of the models
@@ C.18 SeparableKernelClassifier
Appendix D Convergence criteria of variational models
@@
Appendix E Details on the datasets used in this study
@@ E.3 Downscaled MNIST
E.4 Hidden manifold model
@@ E.6 Hyperplanes and parity
Appendix F Collection of detailed results